The length of one side of a square is 13 feet. What is the length, to the nearest foot, of a diagonal of the square
Answer by nerdybill(7384)

The length of one side of a square is 13 feet. What is the length, to the nearest foot, of a diagonal of the square
.
Apply Pythagorean theorem:
Let d = diagonal
then
d^2 = 13^2 + 13^2
d^2 = 169 + 169
d^2 = 338
d = 18.38
rounded to nearest foot
d = 18 feet